the numberical value if x=3 and y=9.
3(x+7y)=

Found 2 solutions by unlockmath, rapaljer:
Answer by unlockmath(1688)

Hello,
Plug these numbers into 3(x+7y)=
3(3+63)=
9+ 189= 198
Make sense?
RJ
www.math-unlock.com

Answer by rapaljer(4671)

How does that Glade commercial go: "Plug it in, plug it in!"

x=3 and y=9
3(x+7y)
3(3+7*9)

Now you have to know the order of operations to know what to calculate first. The first step is to do what is inside the parentheses, and multiplication comes befoe addition:
3(3+63)

Next, add within the parentheses:
3(66)
198

For additional explanation on topics like this, please see my own website. To find my website, click on my tutor name "rapaljer" anywhere in algebra.com. Click on the link there that will take you to the Homepage of my website, which is a subwebsite at Seminole State College of Florida. From my "Homepage", look for the link "Basic, Intermediate, and College Alegebra: One Step at a Time." Click on "Basic Algebra", and look in "Chapter 1" for several topics on Signed Numbers, Order of Operations, and Variables. Here you will find my own complete non-traditional, user-friendly explanation of the topics that my own students found MUCH easier to understand than the traditional textbooks! Many of my exercises here are explained in great detail in the MATH IN LIVING COLOR pages that go with this curriculum. Best of all, it's all FREE!

I have explanations on nearly every topic in algebra and many from geometry as well, and many videos from my younger days as well. Everything is FREE, and I don't even have advertising on the pages!! If you don't need help with math, then visit my website to see the JOKE PAGE--all clean jokes, safe for the little ears in your family!!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida